6 pointsl Suppose that the probability of successfully connecting to the RowanSecure wi-fi network (p) is 0.88 for any individual attempt.
Answer #1

Solution:-

The probability that fewer than six of these attempts will be successful is 0.0037.

p = 0.88, n = 10, x = 6

By applying binomial distribution:-

P(x,n) = nCx*px*(1-p)(n-x)

P(x < 6) = P(x = 1) + P(x = 2) + P(x = 3) + P(x = 4) + P(x = 5)

P(x < 6) = 0.0037